Solid PbI2 was added to a 0.030 M NaI solution. Calculate the molar concentration of lead ion in this solution. Ksp = 7.9 x 10-9 (at 25ºC) (Please show work)

PbI2(s) <=> Pb2+ + 2I-
Answer #1

Answer

8.77×10-6M

Explanation

PbI2(s) <------> Pb2+(aq) + 2I-(aq)

Ksp = [Pb2+] [ I-]2 = 7.9×10-9

Initial concentration

[Pb2+] = 0

[I-] = 0.030

change in concentration

[Pb2+ ] = + x

[I-] = +2x

Equillibrium concentration

[Pb2+] = x

[I-] = 0.030 + 2x

so,

x (0.030 + 2x)2 = 7.9 ×10-9

solving for x

x = 8.77×10-6

Therefore

molar concentration of Pb2+ = 8.77×10-6M